OpenStudy (anonymous):

Quadratic Formula 2x²-7x-13=-10

OpenStudy (anonymous):

2x\[^2\] -7x -13+10=0 2x\[^2\]-7x-3 =0 insert in a formula -b\[\pm \sqrt{b ^{2}-41c}/2a\]

OpenStudy (anonymous):

I insert -7x-3=0 into the formula ?

OpenStudy (anonymous):

2x^2 -7x -3 =0 into the formula a=2 b=-7 c= -3